VEGAS Team Meeting

Dial in Number: 877-937-8961

Passcode: 9445091#

Attendees: Joe, Mark, Richard, Ray, Randy, Jason, Dan, Sam, Mark, Paul, Anish, MArty

Agenda:

Whole group:

  1. Status of BOF files (should all be complete)

GB only

  1. Status of Mode 10-19 firmware fix
  2. Review of recent commissioning observations
  3. Enable real-time streaming
  4. Review of problems from https://safe.nrao.edu/wiki/bin/view/CICADA/CICADAGreenBankSpectrometerCommissioningLogs
  5. Test plans for this week

Meeting notes:

Whole group:

  1. Status of BOF files (should all be complete)
    • Complete. Paul has calibrated the BOF file, should be good to go. This is Mode 2. Now needs astronomical testing

GB only

  1. Status of Mode 10-19 firmware fix
    • Randy has sent a new version to Joe. Joe to fix this week.
  2. Review of recent commissioning observations
    • Generally went well. Some issues with Tsys; may be real astronomical variations.
  3. Enable real-time streaming
    • Enabled. Should be good to go.
  4. Review of problems from https://safe.nrao.edu/wiki/bin/view/CICADA/CICADAGreenBankSpectrometerCommissioningLogs
    • Discussed in detail
    • Main discussion was around spikes in center channel. Dan to discuss further with Glenn and Hong. Richard to write MR for these to be patched in the filler.
  5. Test plans for this week
    • Unscheduled request(s)
      • None
    • Tuesday
      • GBTDATA replacement
      • BOF loading problem investigation
    • Wednesday
      • None at this time
    • Thursday
      • IF power level variations
    • Monday (7/21)
      • None at this time

Other

  1. A 'final' draft of the blanking memo will be circulated by Amanda [Done?]
  2. If we see traffic spikes we could plan to meet additionally on Wednesdays
  3. Please make sure to properly code your timesheets for the current time period and going forward.
    • Software engineers should be using a GBT ops software CM&E account - Account 893252
    • Electronic engineers should be using a GBT ops electronics CM&E account - Account 432510
    • Commissioning activities should be getting charged to hardware support and testing - Account 432510
    • Scientists assisting with shared-risk commissioning should be charging an observer support account
      • 432540 for the most part. Accounts 142020 (GB-based staff) and 142010 (CV-based staff) for when you are helping with overarching observing support (not problem solving, VEGAS specific)

-- MartyBloss - 2014-07-14
Topic revision: r3 - 2015-01-27, RichardPrestage
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback